OpenStudy (anonymous):

What is the distance between points (3, 1) and (0, –5)? Round to the nearest tenth

hartnn (hartnn):

distance between two points is given by \[\sqrt{(x1-x2)^2+(y1-y2)^2}\] does this help?

OpenStudy (anonymous):

Im bot sure how to insert the numbers into the formula. But i tryed and i think its c. Answer choices: A. 11.0 B.19.4 C.8.7 D.6.7

hartnn (hartnn):

\[\sqrt{(3-0)^2+(1-(-5))^2}=\sqrt{9+36}=\sqrt{45}=6.7082\] ok?

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome :)